Dividends Per Share Lightfoot Inc., a software development firm, has stock outstanding as follows: 15,000 shares of cumulative preferred 2% stock, $20 par, and 19,000 shares of $125 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $2,250; second year, $3,750; third year, $28,340; fourth year, $61,480. Calculate the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $.15 $.25 $ $.40
Common stock (dividend per share) $0 $0 $ $2.92

Can't figure out how to get the third year. Thank you.
Answer #1

Number of preferred share = 15,000

Par value per preferred share = $20

Preferred dividend rate = 2%

Dividend per preferred share = Par value per preferred share x Preferred dividend rate

= 20 x 2%

= $0.40

Annual preferred dividends= Number of preferred share x Dividend per preferred share

= 15,000 x 0.40

= $6,000

Total Cash Dividend Paid Paid to Preferred Paid to Common Dividends in Arrears at year - end
Year 1 2,250 2,250 0 3,750
Year 2 3,750 3,750 0 6,000
Year 3 28,340 12,000 16,340 0
Year 4 61,480 6,000 55,480 0
Totals $95,820 $24,000 $71,820 $9,750
1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) 2,250/15,000 = $0.15 3,750/15,000 = $0.25 12,000/15,000 = $0.80 6,000/15,000 = $0.40
Common stock (dividend per share) $0 $0 16,340/19,000 = $0.86 55,480/19,000 = $2.92

Kindly comment if you need further assistance. Thanks‼!